Looking at Andrew's balance sheet, suppose they want to increase its leverage to 3.0 by issuing bonds to purchase plant and equipment. How much additional plant and equipment (in 000's) could it buy?

Group of answer choices

$143,826

$30,026

$67,959

$47,601

Balance Sheet Survey Cash Accounts Receivable Inventory Total Current Assets Plant and equipment Accumulated Depreciation Totnone of the above
Answer #1

answer.

leverage = total assets / total equity

therefore if we the current leverage is

total assets = $96225

total equity = $47942

thus current leverage is 96225 / 47942 = 2

now if we increase the leverage by issuing bonds to 3

there is no change to equity, thus keeping equity constant,

total assets* = equity * leverage*

total assets* = 47942 * 3 = $143826

if we see total assets have increased as new plant and machinery is added to the total assets

change is total assets = 143826 - 47942 = $47601

thus new plant and machinery is bought for $47601 by issuing bonds

thus option d is correct